# All Models Are Miscalibrated, But Some Less So: Comparing Calibration with Conditional Mean Operators

### v1

@@ 3 Conditional Kernel Calibration Error
Recall our notion of calibration given by Definition 2.2 where we compare the difference between probability distributions $P(Y|Q_{X})$ and $Q_{X}$. By defining a random variable $Z_{X}\sim Q_{X}$ obtained from the predictive model, we note that $Q_{X}=P(Z_{X}|Q_{X})$ and hence, a probabilistic predictive model $Q$ is calibrated if and only if $P(Y|Q_{X})=P(Z_{X}|Q_{X})$. Since we are now comparing differences between conditional probability distributions, we apply the CMMD and define the conditional kernel calibration error (CKCE) with respect to kernels $k:\mathcal{P}\times\mathcal{P}\to\mathbb{R}$ and $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ as
@@
A CKCE of 0 implies the conditional distributions are equal almost everywhere, and thus strong calibration of $Q_{X}$ holds. We propose to use the CKCE as a measure of calibration error for relative comparisons, which allows to construct a ranking of a set of candidate models ordered by their calibration.
@@ 3.1 Empirical evaluation
The CKCE can be computed by using empirical estimates of the CMOs as in equation (4). Given a set of $n$ calibration samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ from the joint distribution $P(X,Y)$, we get the predicted conditional distribution of $Y$ given feature $x_{i}$ under our model $Q$ by $q_{i}=f(x_{i})$. Then $\{y_{i},q_{i}\}_{i=1}^{n}$ are samples from $P(Y,Q_{X})$, and the CMO $\mathcal{C}_{Y|Q_{X}}$ can be estimated by
$\widehat{\mathcal{C}}_{Y|Q_{X}}=\Psi_{Y}(K_{QQ}+\lambda nI_{n})^{-1}\Phi_{Q}^{*}$
where $\Psi_{Y}=[\psi(y_{1}),\dots,\psi(y_{n})]$ and $\Phi_{Q}=[\phi(q_{1}),\dots,\phi(q_{n})]$, $K_{QQ}=\Phi_{Q}^{*}\Phi_{Q}$ is the Gram matrix with entries $[K_{QQ}]_{ij}=k(q_{i},q_{j})$, $I_{n}$ is the $n\times n$ identity matrix and $\lambda$ is a regularisation parameter.
The CMO $\mathcal{C}_{Z_{X}|Q_{X}}$ can be estimated in a similar way. However, rather than sampling $Z_{x_{i}}$ from model outputs $Q_{x_{i}}$ to compute the feature matrix $\Psi_{Z_{X}}$, we can replace each of the feature embeddings $\psi(Z_{x_{i}})$ with their expected value, that is, the kernel mean embedding $\mu_{Q_{x_{i}}}$. This reduces variance of the CMO estimator [^1] A similar strategy was used in Chau et al., (2021) and termed Conditional Mean Shrinkage Operator.. Hence,
$\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}=M_{Q}(K_{QQ}+\lambda nI_{n})^{-1}\Phi_{Q}^{*}$
where $M_{Q}=[\mu_{Q_{x_{1}}},\dots,\mu_{Q_{x_{n}}}]$ is a matrix of mean embeddings of the predictive distributions.
Using the CMO estimators, the empirical evaluation of the CKCE is given by
$\widehat{\text{CKCE}}_{k,\ell}=\|\widehat{\mathcal{C}}_{Y|Q_{X}}-\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}\|_{HS}^{2}.$
Since the squared Hilbert-Schmidt norm of an operator can be expressed as $\|\mathcal{C}\|_{HS}^{2}=Tr(\mathcal{C}^{*}\mathcal{C})$, an equivalent formulation for the CKCE estimator is
$\widehat{\text{CKCE}}_{k,\ell}=Tr\Big{(}(K_{QQ}+\lambda nI_{n})^{-1}[M_{Q}^{*}M_{Q}+L_{YY}\\ -M_{Q}^{*}\Psi_{Y}-\Psi_{Y}^{*}M_{Q}](K_{QQ}+\lambda nI_{n})^{-1}K_{QQ}\Big{)}$
where $L_{YY}=\Psi_{Y}^{*}\Psi_{Y}$ is the Gram matrix with entries $[L_{YY}]_{ij}=\ell(y_{i},y_{j})$.
Although estimating the CKCE involves finding the inverse of a potentially large matrix, which is computationally expensive, this can be overcome by a suite of well established large-scale kernel approximations such as random Fourier features (Rahimi and Recht,, 2007). Alternatively, simple low-rank approximations of the Gram matrix $K_{QQ}$ such as incomplete Cholesky factorization (Fine and Scheinberg,, 2002) can be used.
@@ 3.2 Choice of kernel
Evaluating the CKCE requires choosing two kernels, one on the space of labels $\mathcal{Y}$ and another on the space of probability distributions $\mathcal{P}$. We focus on selecting suitable kernels for probabilistic classifiers, where $\mathcal{Y}=\{1,\ldots,m\}$ and $\mathcal{P}=\Delta^{m-1}$. A reasonable choice for the former is the Kronecker kernel $\ell(y,y^{\prime})=\mathbf{1}\{y=y^{\prime}\}$. In this case $\mathcal{H}_{\ell}\cong\mathbb{R}^{m}$ with feature maps $\psi(j)=e_{j}$ as the canonical basis. Then, for a conditional distribution $P(Y|Q_{X})$ described by $\pi:\Delta^{m-1}\to\Delta^{m-1}$, that is, $\pi(q)=P(Y|Q_{X}=q)$, we have
$\mu_{Y|Q_{X}=q}=\mathbb{E}[e_{Y}|Q_{X}=q]=\sum_{j=1}^{m}\pi_{j}(q)e_{j}=\pi(q)$
which is understood as a vector in $\mathbb{R}^{m}$. Thus, the conditional mean embedding is some mapping (potentially nonlinear) of the probability vector $q$. On the other hand, $P(Z_{X}|Q_{X}=q)=q$, and so the conditional mean embedding is simply $\mu_{Z_{X}|Q_{X}=q}=q$.
@@
$k(p,q)=p^{\top}q+\exp\left(-\frac{1}{2\gamma^{2}}\|p-q\|^{2}\right)$
where $\gamma\in\mathbb{R}$ is the bandwidth. An additional benefit of such a kernel is that random Fourier features can easily be used to approximate it, allowing each CMO to be estimated in the primal by applying the Woodbury identity, which speeds up computation for large $n$. Further details on implementation are provided in Appendix B.
The CKCE can likewise be applied for regression problems. For real valued targets, many commonly used kernels such as the Gaussian or Matérn kernels could be chosen. However, note that in those cases we would need to estimate mean embeddings of predictive distributions, which may require sampling $Z_{X}$ from $Q_{X}$. For the space of probability distributions, one could use kernels of the form $k(p,p^{\prime})=\exp(-\gamma D_{\mathcal{P}}(p,p^{\prime}))$ where $D_{\mathcal{P}}$ is an appropriate metric on $\mathcal{P}$. When $D_{\mathcal{P}}$ is the squared MMD, one can in certain cases obtain a universal kernel on probability measures (Christmann and Steinwart,, 2010).
3.3 Beyond global calibration
Definition 2.2 stipulates a global property of a probabilistic model, requiring that it is calibrated on average over all inputs. However, in certain application it may be desirable to check whether model outputs are calibrated for certain groups or even individuals Pleiss et al., (2017); Zhao et al., (2020). This can be cast as conditioning on an extra variable in the LHS of equation (3). For example, let $A$ be a random variable representing the group membership of $X$, which could be a protected attribute such as a person’s race or gender. Then a model $Q$ is calibrated for group $A$ if $P(Y|Q_{X},A)=Q_{X}$. Such extensions of calibration have great importance in the study of algorithmic fairness Chouldechova, (2017); Pleiss et al., (2017). By defining a kernel $k$ on the product space $\mathcal{P}\times\mathcal{A}$, one can similarly use the CKCE to measure the group calibration of a probabilistic model. More generally, the requirement $P(Y|Q_{X},A)=Q_{X}$ could range from $A=\emptyset$, which returns global calibration given by Definition 2.2, to $A=X$, which gives individual calibration Zhao et al., (2020), and everything in between.
Figure 2: Proportion of trials where the estimated calibration error of the calibrated model is less than that of the miscalibrated model for calibration sets of different sizes, with error bars indicating the Wilson score interval.
@@ 4.1 Proper scoring rules
Historically, proper scoring rules such as the Brier score have been used to determine calibration (Brier,, 1950). Other common proper scoring rules are the negative log likelihood and cross entropy, which are both frequently used as loss functions when training modern machine learning models. Since these loss functions are minimised by the true conditional distribution, they can give a well calibrated model with enough training data. The reason behind this is that proper scoring rules can be decomposed into components of calibration and sharpness (Kull and Flach,, 2015). However, complex models can achieve low loss as a result of high accuracy rather than good calibration, meaning these metrics are of limited use for measuring calibration error directly. Furthermore, Merkle and Steyvers, (2013) showed that different scoring rules can rank the same models differently, hampering their use for relative comparison of calibration.
@@ 4.2 Expected Calibration Error
A widely used measure of calibration for probabilistic models is the expected calibration error (ECE) (Naeini et al.,, 2015). Although in its usual formulation ECE measures confidence calibration of probabilistic classifiers, it can be generalised to strong calibration as given by Definition 2.2. Let $d:\mathcal{P}\times\mathcal{P}\to\mathbb{R}_{\geq 0}$ be a metric on the space of probability measures. Then the ECE of a probabilistic model $Q$ with respect to $d$ is defined as
@@
The main difficulty in computing the ECE comes from estimating $P(Y|Q_{X})$. For probabilistic classifiers, an approach suggested by Vaicenavicius et al., (2019) is to partition the probability simplex and use histogram regression, which we implemented in our experiments. The simplest choice is uniform binning, although this causes the number of bins to increase exponentially with the number of classes, meaning care must be taken to ensure enough samples per bin. Alternative binning schemes may also be used, but as demonstrated by Vaicenavicius et al., (2019), the ECE estimator becomes highly sensitive to the chosen scheme further emphasising the need for an alternative calibration metric.
A special case of ECE is the average distance between conditional mean embeddings, which is achieved by choosing $d$ to be the squared MMD. This approach was further explored by Chatterjee et al., (2024), but their method requires taking samples $Z_{X}$ from $Q_{X}$ which can be difficult to marginalise, adding extra variance to the estimator.
@@ 4.3 Joint Kernel Calibration Error
An alternative kernel based measure of calibration considers the discrepancy between joint distributions. As before, we define the random variable $Z_{X}\sim Q_{X}$. Then the joint distributions $P(Y,Q_{X})$ and $P(Z_{X},Q_{X})$ are equal if and only if $P(Y|Q_{X})=Q_{X}$ almost surely, that is, the probabilistic predictive model $Q$ is calibrated (Widmann et al.,, 2021). Let $\kappa:(\mathcal{Y}\times\mathcal{P})\times(\mathcal{Y}\times\mathcal{P})\to\mathbb{R}$ be a measurable kernel with associated RKHS $\mathcal{H}$. Then, the joint kernel calibration error (JKCE) with respect to kernel $\kappa$ is defined as
$\text{JKCE}_{\kappa}=\|\mu_{YQ_{X}}-\mu_{Z_{X}Q_{X}}\|_{\mathcal{H}}^{2}.$ (7)
The JKCE was named the squared kernel calibration error (SKCE) by Widmann et al., (2021) and corresponds to the MMD between the joint distributions $P(Y,Q_{X})$ and $P(Z_{X},Q_{X})$. As described in Section 2.2, for a characteristic kernel $\kappa$ the JKCE equals 0 if and only if the distributions $P(Y,Q_{X})$ and $P(Z_{X},Q_{X})$ are equal, i.e., $Q$ is calibrated. Given kernels $\kappa_{\mathcal{Y}}$ on the space of labels and $\kappa_{\mathcal{P}}$ on the space of probability measures, a kernel on the product domain $\mathcal{Y}\times\mathcal{P}$ can be constructed by their tensor product, i.e., $\kappa=\kappa_{\mathcal{Y}}\otimes\kappa_{\mathcal{P}}$. If $\kappa_{\mathcal{Y}}$ and $\kappa_{\mathcal{P}}$ are universal kernels on locally compact Polish spaces, then $\kappa=\kappa_{\mathcal{Y}}\otimes\kappa_{\mathcal{P}}$ is characteristic (Szabó and Sriperumbudur,, 2018, Theorem 5). Since the kernels $k$ and $\ell$ from Section 3.2 are universal, setting $\kappa=\ell\otimes k$ produces a characteristic kernel. We will use this kernel to evaluate JKCE in our experiments in Section 5 for a fair comparison with CKCE.
Given a set of calibration samples, it is possible to construct an unbiased estimator $\widehat{\text{JKCE}}_{\kappa}$ (Widmann et al.,, 2021) and use it to test the null hypothesis of strong calibration. However, the drawback is that when applied to two candidate models $Q^{1}$ and $Q^{2}$, the estimator will be sensitive to the difference between the marginal distributions of $Q^{1}_{X}$ and $Q^{2}_{X}$, which is not itself relevant for calibration.
Table 1: Accuracy, cross entropy and calibration error (using different estimators, scaled to be roughly the same order of magnitude) of neural network architectures and the marginal probability model on CIFAR-10 and ImageNet datasets.

### v2

@@ 3 Conditional Kernel Calibration Error
Recall our notion of calibration error described in Section 2.1 where we compare the difference between probability distributions $P(Y|Q_{X})$ and $Q_{X}$. By defining a random variable $Z_{X}\sim Q_{X}$ obtained from the predictive model, we note that $Q_{X}=P(Z_{X}|Q_{X})$ and hence, a probabilistic predictive model is calibrated if and only if $P(Y|Q_{X})=P(Z_{X}|Q_{X})$ almost surely. Since we are now comparing differences between conditional probability distributions, we apply the CMMD and define the conditional kernel calibration error (CKCE) with respect to kernels $k:\Delta^{m-1}\times\Delta^{m-1}\to\mathbb{R}$ and $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ as
@@
A CKCE of 0 implies the conditional distributions are equal almost everywhere, and thus strong calibration holds. Since the CKCE is not impacted by the marginal distribution of $Q_{X}$, it provides a measure of calibration error for relative comparisons of candidate models.
@@ 3.1 Empirical Evaluation
The CKCE can be computed by using empirical estimates of the CMOs as in equation (4). Given a set of $n$ calibration samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ from the joint distribution $P(X,Y)$, we compute prediction $q_{i}=f(x_{i})$ for each feature using the classifier. Then $\{(y_{i},q_{i})\}_{i=1}^{n}$ are samples from $P(Y,Q_{X})$, and the CMO $\mathcal{C}_{Y|Q_{X}}$ is estimated by $\widehat{\mathcal{C}}_{Y|Q_{X}}=\Psi_{Y}(K_{QQ}+\lambda nI_{n})^{-1}\Phi_{Q}^{*}$.
The CMO $\mathcal{C}_{Z_{X}|Q_{X}}$ can be estimated in a similar way. However, rather than sampling $Z_{x_{i}}$ from model outputs $q_{i}$ to construct the feature matrix $\Psi_{Z_{X}}$, we can replace each of the feature embeddings $\psi(Z_{x_{i}})$ with their expected value, that is, the kernel mean embedding $\mu_{Z_{x_{i}}}$. This reduces variance of the CMO estimator [^1] A similar strategy was used in [ChaBouSej2021] and termed Conditional Mean Shrinkage Operator.. Hence, $\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}=M_{Z}(K_{QQ}+\lambda nI_{n})^{-1}\Phi_{Q}^{*}$ where $M_{Z}=[\mu_{Z_{x_{1}}},\dots,\mu_{Z_{x_{n}}}]$ is a matrix of mean embeddings of the predictive distributions. Using the CMO estimators, the empirical evaluation of the CKCE is given by
$\widehat{\text{CKCE}}_{k,\ell}$ $=\|\widehat{\mathcal{C}}_{Y|Q_{X}}-\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}\|_{HS}^{2}$ (7)
$=Tr\Big(W[M_{Z}^{*}M_{Z}+L_{YY}-M_{Z}^{*}\Psi_{Y}-\Psi_{Y}^{*}M_{Z}]WK_{QQ}\Big)$
where $W=(K_{QQ}+\lambda nI_{n})^{-1}$, $L_{YY}=\Psi_{Y}^{*}\Psi_{Y}$ is the Gram matrix with entries $L_{YY}]_{ij}=\ell(y_{i},y_{j})$, and we have applied the fact that the squared Hilbert-Schmidt norm of an operator can be expressed as $\|\mathcal{C}\|_{HS}^{2}=Tr(\mathcal{C}^{*}\mathcal{C})$.
Although estimating the CKCE involves finding the inverse of a potentially large matrix, which is computationally expensive, this can be overcome by a suite of well established large-scale kernel approximations such as random Fourier features [rahimiRandomFeaturesLargescale2007], which we explore in Section 3.3. Alternatively, simple low-rank approximations of the Gram matrix $K_{QQ}$ such as incomplete Cholesky factorization [Fine_2002] can be used for further computational efficiency.
@@ 3.2 Choice of Kernel
Evaluating the CKCE requires choosing two kernels, one on the space of labels $\mathcal{Y}$ and another on the probability simplex $\Delta^{m-1}$. A reasonable choice for the former is the Kronecker kernel $\ell(y,y^{\prime})=\mathbf{1}\{y=y^{\prime}\}$. In this case $\mathcal{H}_{\ell}\cong\mathbb{R}^{m}$ with feature maps $\psi(j)=e_{j}$ as the canonical basis. Then, for a conditional distribution $P(Y|Q_{X})$ described by $\pi(q)=P(Y|Q_{X}=q)$, we have $\mu_{Y|Q_{X}=q}=\mathbb{E}[e_{Y}|Q_{X}=q]=\sum_{j=1}^{m}\pi_{j}(q)e_{j}=\pi(q)$, which is understood as a vector in $\mathbb{R}^{m}$. Thus, the conditional mean embedding is some mapping (potentially nonlinear) of the probability vector $q$. On the other hand, $P(Z_{X}|Q_{X}=q)=q$, and so the conditional mean embedding is simply $\mu_{Z_{X}|Q_{X}=q}=q$.
@@
$k(p,q)=p^{\top}q+\exp\left(-\frac{1}{2\gamma^{2}}\|p-q\|^{2}\right)$ (8)
where $\gamma\in\mathbb{R}$ is the bandwidth. Such a combination of the linear and Gaussian kernels ensures that the CMO assumptions are reasonably satisfied, and establishes robustness of the CKCE metric which we demonstrate experimentally in Section 5.3.
3.3 Random Fourier Feature Implementation
An additional benefit of kernel (8) is that it can be easily approximated by random Fourier features. This allows each CMO to be estimated in the primal by applying the Woodbury identity, which speeds up computation for large $n$ without sacrificing convergence properties [liUnifiedAnalysisRandom2021]. Since adding kernels is equivalent to concatenating features, an approximation of the canonical feature map for kernel $k$ is
$\hat{\phi}(p)=\left[p^{\top},\frac{1}{\sqrt{D}}\cos(w_{1}^{\top}p),\frac{1}{\sqrt{D}}\sin(w_{1}^{\top}p),\dots,\frac{1}{\sqrt{D}}\cos(w_{D}^{\top}p),\frac{1}{\sqrt{D}}\sin(w_{D}^{\top}p)\right]^{\top}$
where $p\in\Delta^{m-1}$ and $w_{1},\dots,w_{D}$ are $D$ independent samples from the multivariate normal distribution $\mathcal{N}_{m}(0,\gamma^{-2}I)$. One can show that $\mathbb{E}[\hat{\phi}(p)^{\top}\hat{\phi}(q)]=k(p,q)$ and $\text{Var}[\hat{\phi}(p)^{\top}\hat{\phi}(q)]=O(D^{-1})$ [rahimiRandomFeaturesLargescale2007]. In our experiments, we choose $D=100$.
Given a calibration set $\{(y_{i},q_{i})\}_{i=1}^{n}$, where each $q_{i}$ is the probability vector predicted by our classifier for a feature $x_{i}$, the CMO estimators are
$\widehat{\mathcal{C}}_{Y|Q_{X}}=\Psi_{Y}\widehat{\Phi}_{Q}^{*}(\widehat{\Phi}_{Q}\widehat{\Phi}_{Q}^{*}+\lambda nI_{\tilde{m}})^{-1},\qquad\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}=M_{Z}\widehat{\Phi}_{Q}^{*}(\widehat{\Phi}_{Q}\widehat{\Phi}_{Q}^{*}+\lambda nI_{\tilde{m}})^{-1}$ (9)
where $\widehat{\Phi}_{Q}=[\hat{\phi}(q_{1})\dots\hat{\phi}(q_{n})]$ and $I_{\tilde{m}}$ is the identity matrix with $\tilde{m}=m+2D$. The resulting CKCE estimator is $\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}-\widehat{\mathcal{C}}_{Y|Q_{X}}||_{F}^{2}$ where we have replaced the Hilbert-Schmidt norm with the Frobenius norm as we are now working with finite dimensional matrices rather than operators. As the main computational complexity comes from calculating the inverse, we have thus reduced the computational cost from $O(n^{3})$ to $O(\tilde{m}^{3})$ with $\tilde{m}\ll n$.
@@ 4.1 Proper Scoring Rules
Historically, proper scoring rules such as the Brier score have been used to determine calibration [Brier_1950]. Other common proper scoring rules are the negative log likelihood and cross entropy, which are both frequently used as loss functions when training modern machine learning models. Since these loss functions are minimised by the true conditional distribution, they can give a well calibrated model with enough training data. However, complex models can achieve low loss as a result of high accuracy rather than good calibration, meaning these metrics are of limited use for measuring calibration error directly. Furthermore, [Merkle_2013] showed that different scoring rules can rank the same models differently, hampering their use for relative comparison of calibration.
@@ 4.2 Expected Calibration Error
A widely used measure of calibration for probabilistic models is the expected calibration error (ECE) [naeini_2015]. Although in its usual formulation ECE measures confidence calibration of probabilistic classifiers, it can be generalised to strong calibration as given by Definition 1. Let $d$ be a metric on $\Delta^{m-1}$. Then the ECE of a probabilistic model with respect to $d$ is defined as
@@
The main difficulty in computing the ECE comes from estimating $P(Y|Q_{X})$. An approach suggested by [vaicenavicius_2019] is to partition the probability simplex and use histogram regression, which we implemented in our experiments. The simplest choice is uniform binning, although this causes the number of bins to increase exponentially with the number of classes, meaning care must be taken to ensure enough samples per bin. Other binning schemes may also be used, but as demonstrated by [vaicenavicius_2019], the ECE estimator becomes highly sensitive to the chosen scheme further emphasising the need for an alternative calibration metric. A special case of ECE is the average distance between conditional mean embeddings, which is achieved by choosing $d$ to be the squared MMD. This approach was further explored by [Chatterjee_2024], but their method requires taking samples $Z_{X}$ from $Q_{X}$ which can be difficult to marginalise, adding extra variance to the estimator.
@@ 4.3 Joint Kernel Calibration Error
An alternative kernel based measure of calibration considers the discrepancy between joint distributions. As before, we define the random variable $Z_{X}\sim Q_{X}$. Then the joint distributions $P(Y,Q_{X})$ and $P(Z_{X},Q_{X})$ are equal if and only if $P(Y|Q_{X})=Q_{X}$ almost surely, that is, the probabilistic predictive model is calibrated [Widmann_2021]. Let $\kappa:(\mathcal{Y}\times\Delta^{m-1})\times(\mathcal{Y}\times\Delta^{m-1})\to\mathbb{R}$ be a measurable kernel with associated RKHS $\mathcal{H}_{\kappa}$. Then, the joint kernel calibration error (JKCE) with respect to kernel $\kappa$ is defined as
$\text{JKCE}_{\kappa}=\|\mu_{YQ_{X}}-\mu_{Z_{X}Q_{X}}\|_{\mathcal{H_{\kappa}}}^{2}.$ (11)
The JKCE corresponds to the MMD between the joint distributions $P(Y,Q_{X})$ and $P(Z_{X},Q_{X})$. As described in Section 2.2, for a characteristic kernel $\kappa$ the JKCE is 0 if and only if the model is calibrated. Given kernels $\kappa_{\mathcal{Y}}$ on the space of labels and $\kappa_{\Delta}$ on the probability simplex, a kernel on the product domain $\mathcal{Y}\times\Delta^{m-1}$ can be constructed by their tensor product. If $\kappa_{\mathcal{Y}}$ and $\kappa_{\Delta}$ are universal kernels on locally compact Polish spaces, then $\kappa=\kappa_{\mathcal{Y}}\otimes\kappa_{\mathcal{P}}$ is characteristic [Szabo_2018, Theorem 5]. Since the kernels $k$ and $\ell$ from Section 3.2 are universal, setting $\kappa=\ell\otimes k$ produces a characteristic kernel. We will use this kernel to evaluate JKCE in our experiments in Section 5 for a fair comparison with CKCE.
5 Experimental Results
In our experiments, we assess the ability of the CKCE to distinguish between models through a relative comparison of calibration error. We compare its performance with ECE and JKCE. In the following experiments, the kernel in equation (8) is used on the input variable for both the CKCE and JKCE, and the bandwidth parameter $\gamma$ is chosen via the median heuristic [Gretton_2012]. The ECE is estimated by uniform binning of the simplex, and we take $d$ to be the $L^{1}$ norm as is common in the literature [Guo_2017, vaicenavicius_2019].
Table 1: Performance metrics and calibration error of neural network architectures and the marginal probability model on the ImageNet dataset.